The distance between the eyepiece and the objective lens in a certain compound microscope is 23.0cm. The focal length of the eyepiece is 2.50cm and that of the objective is 0.400cm. What is the overall magnification of the microscope?


1
Expert's answer
2021-03-17T19:53:38-0400

We have the following formula to solve this problem

"L = 23.0cm \space\space\space\space f_o=0.400cm \space\space\space\space D_0 = 25cm \space\space\space\space f_e = 2.50cm"

"M = - \large(\frac{L}{f_o})(\frac{D_o}{f_e}) = - (\frac{23.0cm}{0.400cm})(\frac{25.0cm}{2.50cm})" "= - 575"